OpenStudy (anonymous):

If I have an 85% in my class and my final is worth 20% of my final grade, what do I need to get on my final to get at least an 80% in the class?

OpenStudy (anonymous):

solve \[.8\times 85 + .2\times x \geq 80\] for x

OpenStudy (anonymous):

i.e. 80 percent of your grade is 85, twenty percent is the final (which you don't know yet) and you want that larger than 80 can you solve this?

OpenStudy (anonymous):

I think so...thanks =)

OpenStudy (anonymous):

answer is > 60

OpenStudy (anonymous):

AWESOME! That's what I got too! Thank you! That puts my mind at ease heading into my final exam.

OpenStudy (anonymous):

good work